OpenStudy (karatechopper):

You make a purchase at a local hardware store, but what you've bought is too big to take home in your car. For a small fee, you arrange to have the hardware store deliver your purchase for you. You pay for your purchase, plus the sales taxes, plus the fee. The taxes are 7.5% and the fee is $20.

OpenStudy (karatechopper):

(i) Write a function t(x) for the total, after taxes, on the purchase amount x. Write another function f(x) for the total, including the delivery fee, on the purchase amount x. (ii) Calculate and interpret ( f o t)(x) and (t o f )(x). Which results in a lower cost to you? (iii) Suppose taxes, by law, are not to be charged on delivery fees. Which composite function must then be used?

OpenStudy (karatechopper):

(i) Write a function t(x) for the total, after taxes, on the purchase amount x. Write another function f(x) for the total, including the delivery fee, on the purchase amount x. (ii) Calculate and interpret ( f o t)(x) and (t o f )(x). Which results in a lower cost to you? (iii) Suppose taxes, by law, are not to be charged on delivery fees. Which composite function must then be used?

ganeshie8 (ganeshie8):

did u understand wat they asking ?

OpenStudy (karatechopper):

Not really.

ganeshie8 (ganeshie8):

basically they want you write two functions : t(x) and f(x) t(x) : represents cost x + tax of 7.5 % f(x) : represents cost x + delovery fee of $20 flat

ganeshie8 (ganeshie8):

once you have the functions, working rest is bit easy..

OpenStudy (karatechopper):

Oh ok. Yeah..ok.. How would you write the functions tho?

ganeshie8 (ganeshie8):

say, the cost is x

OpenStudy (karatechopper):

ok

ganeshie8 (ganeshie8):

t(x) : represents cost x + tax of 7.5 % t(x) = x + 7.5 % of x simplify

OpenStudy (karatechopper):

is there a dif way to write tax of 7.5%?

ganeshie8 (ganeshie8):

7.5% of goods purchased = 7.5/100*x = 0.075x

ganeshie8 (ganeshie8):

overall, t(x) = x + 7.5 % of x = x + 0.075x = x(1+.075) = 1.075x

ganeshie8 (ganeshie8):

t(x) = 1.075x --------------------------(1)

ganeshie8 (ganeshie8):

similarly work the f(x)

OpenStudy (karatechopper):

Where did the 1 come from!?

ganeshie8 (ganeshie8):

t(x) = x + 7.5 % of x = x + 0.075x = x(1+.075) ******************* = 1.075x

ganeshie8 (ganeshie8):

you're asking that 1 ?

OpenStudy (karatechopper):

Yesh.

ganeshie8 (ganeshie8):

look at the line right above that :- x + 0.075x

ganeshie8 (ganeshie8):

ive just factored out x

OpenStudy (karatechopper):

I don't get it..

OpenStudy (karatechopper):

Oh wait..

OpenStudy (karatechopper):

I GET IT!

OpenStudy (karatechopper):

Oh wait.. no I don't. Shouldn't the 2 x's stay?

ganeshie8 (ganeshie8):

:) here it is x + 0.075x ^ ^

ganeshie8 (ganeshie8):

we see that x exists in both terms, so we pull that out

ganeshie8 (ganeshie8):

x(1 + 0.075)

ganeshie8 (ganeshie8):

when we pull out x from first term, 1 stays there

OpenStudy (karatechopper):

Oh ok.

ganeshie8 (ganeshie8):

when we pull out x from second term, 0.075 stays there

OpenStudy (karatechopper):

So then would it become t(x)=1.0755

OpenStudy (karatechopper):

*t(x)=1.075

ganeshie8 (ganeshie8):

who ate x ? ;)

OpenStudy (karatechopper):

But..u pulled it out..and shemurr?!

ganeshie8 (ganeshie8):

x(1 + 0.075) x(1.075) 1.075x

OpenStudy (karatechopper):

Oh wait.. i see..

ganeshie8 (ganeshie8):

we pulled it out , but it dint go anywhere. it stuck there

ganeshie8 (ganeshie8):

okay, once we make peace wid t(x) function, t(x) = 1.075x ---------------------------(1)

ganeshie8 (ganeshie8):

save it, and put the rquation for f(x)... once we have both funcitons, we can work the rest

ganeshie8 (ganeshie8):

*equation

ganeshie8 (ganeshie8):

f(x) : represents cost x + delovery fee of $20 flat

OpenStudy (karatechopper):

So would that be, x+20?

ganeshie8 (ganeshie8):

perfect ! f(x) = x+20 ---------------------------(2)

OpenStudy (karatechopper):

whats that 2 for?

ganeshie8 (ganeshie8):

im just labelling it as number 2, so that we can refer to that later its just for communication between us

OpenStudy (karatechopper):

ah ok

OpenStudy (karatechopper):

So this is about as far as I got on the second step..

ganeshie8 (ganeshie8):

we're done wud part1, ever dealt wid composition functions before ? fog etc...

OpenStudy (karatechopper):

Yeah I have. just word problems are not my thing lol

ganeshie8 (ganeshie8):

yeah who doesnt lol

OpenStudy (karatechopper):

f o t(x)=1.075x + 20 t o f(x)=22.575x

OpenStudy (karatechopper):

lol

ganeshie8 (ganeshie8):

fot(x) looks correct check tof(x) again

OpenStudy (karatechopper):

t o f(x) t(x+20) t(x)=1.075x t(x)=1.075x(x + 20) = 1.075x^2 + 21.5x

OpenStudy (karatechopper):

that?^

ganeshie8 (ganeshie8):

Yup looks good :)

ganeshie8 (ganeshie8):

lets put all equations we go in one place

ganeshie8 (ganeshie8):

t(x) = 1.075x --------------------------(1) f(x) = x+20 ---------------------------(2) (fot)(x) = 1.075x+20 ---------------------(3) (tof)(x) = 1.075x + 21.5 ----------------(4)

ganeshie8 (ganeshie8):

ur calculatioon, above has a mistake. see correct working below :- t o f(x) t(x+20) t(x)=1.075x t(x+20)=1.075(x + 20) = 1.075x + 21.5

OpenStudy (karatechopper):

ohh ok. i gotcha now

OpenStudy (karatechopper):

So now it asks which equation out of the 3rd or 4th is lower

OpenStudy (karatechopper):

in cost

OpenStudy (karatechopper):

which that means the 3rd is lower

ganeshie8 (ganeshie8):

Sure ?

OpenStudy (karatechopper):

yes I am sure

ganeshie8 (ganeshie8):

put aside equations for some time u purchased $100 Is having 7.5% tax first, followed by fee of $20 a better option, or Is paying $20 first, followed by 7.5% tax a better option ?

ganeshie8 (ganeshie8):

Ahh you're right actually i was thinking something else lol

OpenStudy (karatechopper):

Uuhm the tax first right?

OpenStudy (karatechopper):

lol ok

OpenStudy (karatechopper):

yyayyyyy!!!

ganeshie8 (ganeshie8):

haa u got it !!!! always, have tax on low amount... so that we wud have to pay less tax to govt

OpenStudy (karatechopper):

yeah haha

OpenStudy (karatechopper):

Ok now 3rd part

ganeshie8 (ganeshie8):

you can answer that easily. it says, ur govt is bit generous and it wont charge tax on delivery fee, so we must use which composition funciton ? 3 or 4 ?

OpenStudy (karatechopper):

the same right?

OpenStudy (karatechopper):

3rd

ganeshie8 (ganeshie8):

thikn of it like this :- which function shows, tax is not done on delovery fee ?

ganeshie8 (ganeshie8):

yup !

ganeshie8 (ganeshie8):

we wud ude 3rd function, when govt is good, and we dont have to pay tax on delivery fee

OpenStudy (karatechopper):

ohh ok. THanks!

ganeshie8 (ganeshie8):

*use

OpenStudy (karatechopper):

Lol govt good.

ganeshie8 (ganeshie8):

np :) no govt is good i suppose... lol

ganeshie8 (ganeshie8):

we're still paying taxes

OpenStudy (karatechopper):

lolol

OpenStudy (karatechopper):

That world problem would be like a dream come true haha

ganeshie8 (ganeshie8):

lol yes tax free economy is possible oly in books

OpenStudy (karatechopper):

lol